OpenStudy (anonymous):

sec theta = square root of 2. Where theta lies in Quadrant 1. What is tan theta.

OpenStudy (campbell_st):

 so\[1/\cos = \sqrt{2} \] \[\cos = 1/\sqrt{2}\] use pythagoras... but missing side is 1 unit tan = 1

OpenStudy (campbell_st):

oops 1 should be on the x axis,.... and you need to find the opposite side

Directrix (directrix):

1 + tan theta squared = sec theta squared tan theta squared = 2 - 1 tan theta = 1 (positive in Quad I)

OpenStudy (anonymous):

why is the missing side one unit?